OpenStudy (deadshot):

How do I simplify logv6 4x + logv6 5x + logv6 2x?

OpenStudy (deadshot):

\[\log _{6} 4x+\log _{6} 5x+\log _{6} 2x\]

OpenStudy (anonymous):

\[\log_{a}b+\log_{a}c=\log_{a}bc \]

OpenStudy (loser66):

@diongory good for you, I am glad to see your stuff here :)

OpenStudy (anonymous):

@Loser66 thank you